OpenStudy (anonymous):

Lonnie has a gift card worth $600 for a local entertainment store. Movies cost $20 each and newly released video games cost $50 each. In order to use the gift card, Lonnie must purchase at least thirteen items. Which is a combination of movies and newly released video games that Lonnie could purchase using his gift card? 8 movies and 6 video games 14 movie and 7 video games 6 movies and 4 video games 2 movies and 10 video games

OpenStudy (anonymous):

You have to try plotting in the numbers. For example 14 movies and 7 video games: \[14*20 = 280\] \[7*50=350\] Then we add the two: \[280+350 = 630\] So b is not the answer, now you try :)